Assume that a sample is used to estimate a population proportion p. Find the margin of error E that corresponds to the given statistics and confidence level. Round the margin of error to four decimal places.

95% confidence; n = 334, x = 44
Answer #1

solution :

Given that,

n = 334

x = 44

\hat p = x / n = 44/ 334 = 0.132

1 -\hat p = 1 - 0.132 = 0.868

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

Margin of error = E = Z\alpha / 2 * \sqrt((\hat p * (1 - \hat p )) / n)

= 1.96 * (\sqrt((0.132 * 0.868) / 334) = 0.0363